The device parameters for an n-Channel JFET are: Maximum current iDSS = 10mA. Pinch off voltage, Vp = - 4V Calculate the drain current for (a) VGS = 0 (b) VGS - 1.0v (c) VGS = - 4V.


1
Expert's answer
2022-02-01T07:09:01-0500

Drain current= IDSS (1- Vgs/Vp)2

IDSS = 10mA

Vp = - 4V

(a)-Drain current= 10mA (1- 0)2=10mA

(b)-Drain current= 10mA (1-1/4)2= 3.125mA

(c)-Drain current= 10mA (1-4/4)2= 0mA